The speed of a hiker averages 2 km per hour uphill and 6 km per hour downhill.What is his average speed for the entire trip?
Answer by stanbon(75887)

The speed of a hiker averages 2 km per hour uphill and 6 km per hour downhill.What is his average speed for the entire trip?
--------
Let the distance be "x" km.
-------
Uphill DATA:
Distance = x km ; rate = 2 km/h ; time = d/r = x/2 hrs
-----------------
Downhill DATA:
Distance = x km ; rate = 6 km/h ; time = d/r = x/6 hrs
-----------------
Total Trip DATA:
Distance = 2x km ; Time = x/2 + x/6 = 8x/12 = 2x/3 hrs
----------
Total Trip Rate:
d/t = r
(2x)/(2x/3) = 3 km/hr
===================
Cheers,
Stan H.